# AI-enhanced collective intelligence

### v3

@@ Introduction
Much research on collective intelligence has been dedicated to finding a group’s optimum size and structure. Several studies have reported that performance improves with group size through enhanced diversity 17, whereas others suggest the opposite: small groups are more efficient 125. Another study suggests that large teams develop and small teams innovate 126. Larger groups can better utilize diverse perspectives and knowledge to solve complex problems but also tend to experience more coordination problems and communication difficulties 15. The consideration of optimal group size can be related to the dynamics of the group, whether it is static, growing, or diminishing.
In human-AI hybrid teams, group size becomes nuanced with AI integration, especially when AI entities such as Large Language Models (LLMs) are involved. The countability of AI agents, particularly if multiple agents are derived from a single LLM, raises questions about their distinctiveness and individuality in a team context. Unlike humans, AI agents can simultaneously process multiple tasks, blurring the traditional boundaries of group size. Applying past research on human group dynamics to these hybrid systems requires rethinking the notions of individual contribution, team cohesion, and communication. It is important to consider how AI’s unique capabilities and scale impact these dynamics and whether multiple AI agents from a single model represent distinct entities or a collective intelligence resource.
@@
Research shows that network structure, specifically hierarchy and link intensity, can affect information distribution 127, influencing group performance. Another study observed that the presence of structural holes 128 in leaders’ networks and the adoption of core-periphery and hierarchical structures in groups correlated negatively with their performance 129. Incorporating the temporal dimension, adaptive social networks can significantly impact collective intelligence by allowing the network’s structure to evolve based on feedback from its members 130.
Team size moderates the relationship between network structure and performance 131. Studies have found central members can coordinate with other team members more easily in smaller teams 132, whereas in larger teams, communication challenges may arise 133. A recent study shows a more central connected leader in advice-giving networks has a more positive impact on the performance of larger teams, but this effect is reversed in smaller teams 131.
@@
Teams often outperform individuals primarily due to explicit communication and feedback within the team 134. Understanding the communication network in crowds is crucial for effectively designing and managing crowdsourcing tasks 135. Various studies highlight that centralized and decentralized communication patterns can effectively promote team performance, contingent on the nature of the task and the team’s composition 136, 137. In particular, teams handling complex tasks tend to be more productive with decentralized communication networks 138. Conversely, the centralization of communication around socially dominant or less reflective individuals can negatively impact the utilization of expertise and overall team performance 139. Communication patterns, such as equal participation and turn-taking of team members, can positively affect collective intelligence 15.
The integration of AI in human groups can influence the conversation dynamics in different ways, depending on the role of AI. AI chat interventions can improve online political conversations at scale 140, reducing the chance of conflicts. While AI machines might struggle with capturing the subtle and ineffable social expressions that make up the dynamics of human groups, this can be advantageous. In certain contexts, AI can be social catalysts 141 to promote communication where human capacity is limited. Striking the right balance in AI’s role is key to maintaining the natural dynamics of human interaction.
@@
Cognitive processes are foundations for the emergence of collective intelligence 6, encompassing knowledge acquisition, information processing, problem-solving, decision-making, language, perception, memory, attention, and reasoning. Human cognition processes evolve and interact with each other during various tasks and experiences. Humans can represent the mental states of others, including desires, beliefs, and intentions 142, known as the Theory of Mind (ToM). Past research, inspired by ToM, showed that team members’ ability to assess others’ mental state is positively associated with team performance both in face-to-face settings 15 and online 143. Regarding collective understanding within a group, the Shared Mental Models (SMM) theory is helpful for understanding, predicting, and improving performance in human teams 144.
Applying theories such as SMM to human-AI teams is a natural progression in research. These models can enhance team performance by ensuring that human and AI members have a similar understanding of tasks and each other 144, facilitating better prediction of needs and behaviors. Developing ToM in AI also improves coordination and adaptability in human-AI teams 145. This involves AI’s capability to anticipate and respond to new information and behaviors. Moreover, recent studies indicate that incorporating “hot” cognition, influenced by emotional states in AI, can further improve human-machine interactions 146. Understanding the transactive systems 19 of memory, attention, and reasoning within these teams is key to grasping the members’ knowledge and task preferences.
@@
Assistant-type AI systems generally function as technical tools with limited autonomy, designed to complement or augment human abilities and enhance efficiency in task performance. For example, language translation AI can assist in translating text or spoken language, facilitating communication across language barriers. LLMs in education can assist students’ learning by adapting to different roles based on the prompts provided by the students 147. Administrative AI assistants can help schedule meetings, manage emails, and organize tasks. Smart home AI assistants such as Alexa and Siri can help manage smart home devices, play music, provide weather updates, answer questions, and assist with daily routines. AI assistants are technical tools that help humans coordinate tasks efficiently, optimize their decision-making, and personalize their experiences.
@@
AI can function as teammates, working alongside humans with complementary abilities. AI teammates are already used in real-world settings; most employees using AI already see it as a coworker 148. In healthcare, radiologists and AI work together to diagnose from the pictures of the patients. In creative industries, AI collaborators in music, art, and literature, such as Google’s Magenta 149 project, collaborate with musicians and artists to create new compositions or artworks using AI algorithms. Furthermore, AI robots can serve as teammates in various workplaces, a topic we will discuss in more detail later.
@@
As a coach, AI can provide guidance, support, and personalized feedback to individuals seeking to improve their skills and achieve their goals. In team settings, one perspective highlights the AI coach’s ability to provide a comprehensive, global view of the team’s environment, guiding players who have only partial views 150. This approach involves the AI coach strategizing and coordinating, distributing tailored strategies to each team member based on their unique perspectives and roles. Another dimension of AI coaching focuses on assessing and improving teamwork by closely monitoring team members during collaborative tasks 151. Here, the AI coach actively intervenes with timely suggestions based on the team members’ inferred mental model misalignment 151.
@@
AI has been used in working environments with supervision or managerial roles, such as making decisions of hiring, promotion, and reassigning tasks 152. An AI manager may mitigate human biases 153 in hiring and promotions, contributing to a more diverse workforce, which is a key aspect of collective intelligence. Additionally, by analyzing individual strengths and weaknesses and efficiently allocating tasks, an AI manager can optimize team dynamics and workflows, improving overall group performance.
@@
With the advancement of AI in areas such as natural language processing, object recognition, and creative idea generation, there’s a trend to integrate AI with robotics, creating physical entities in the real world. Embodied AI 154 refers to intelligent agents interacting with their environment through a physical body. These AI-enabled machines, embodied in robotic arms or autonomous vehicles, can augment the capabilities of human workers. They are equipped with advanced sensors, motors, and actuators, enabling recognition of people and objects and safe collaboration with humans in diverse settings, including factories, warehouses, and laboratories. Furthermore, advanced AI-robotics integration allows robots to recognize and respond to human speech and gestures, impacting conversational dynamics within mixed human-robot environments. Research has shown that a vulnerable robot’s social behavior positively shapes the conversation dynamics among human participants in a human-robot group 155. In scenarios where conflicts emerge within a human group or in groups of strangers, a humorous robot can potentially serve as an ice breaker to intervene and resolve the awkward situation 156, demonstrating the potential of robots to influence the nature of human-human interactions notably.
@@
In addition to investigating how humans perceive individual AI, previous research suggests the necessity of additional work to explore human perceptions of AI in collaborative team environments 157. For human-AI teams to succeed, human team members must be receptive to their new AI counterparts 158. People’s pre-existing attitudes toward AI were found to be significantly related to their willingness to be involved in human-AI teams 159. Research finds that two aspects of social perception, warmth, and competence, are critical predictors of human receptivity to a new AI teammate. Psychological acceptance is positively related to perceived human-AI team viability 158. Regarding human perception of AI behaviors, one study finds that people cooperate less with benevolent AI agents than with benevolent humans 160.
Humans tend to be more accepting of AI when it makes mistakes. Some studies find that users have a more favorable impression of the imperfect robot than the perfect robot when the robot behaves adequately after making mistakes 161. When both human drivers and AI make errors in cases of human-machine shared-control vehicles, the blame attributed to the machine is reduced 162. Research has found that errors occasionally performed by a humanoid robot can increase its perceived human likeness and likability 163. A recent study shows humans accept their AI teammate’s decision less often when they are deceived about the identity of the AI as another human 164. The nature of the task requested by the robot, e.g., whether its effects are revocable as opposed to irrevocable, significantly impacts participants’ willingness to follow its instructions 165. When taking AI’s advice, humans must know the AI system’s error boundary and decide when to accept or override AI’s recommendation 166.
@@
In AI-assisted decision-making, where the individual strengths of humans and AI optimize the joint decision outcome, the key to success is appropriately calibrating human trust 167. Trust between humans and AI plays a pivotal role in realizing collective intelligence 168. Previous research highlights the necessity of understanding factors that enable or hinder the formation, maintenance, and repair of trust in human-AI collaborations 168. Key factors influencing this trust include the perceived competence, benevolence, and integrity 169 of the AI systems, mirroring the trust dynamics in human relationships.
The level of anthropomorphism of machines can be a predictor of the trust level of the human participants 170. Studies indicate that while anthropomorphic features in AI agents initially create positive impressions, this effect is often short-lived 168. Trust tends to decline, especially when the agent’s capabilities are not clearly presented and its performance fails to meet users’ expectations 171.
Trust in a teammate may evolve over time irrespective of team performance 172. Knowing when to trust the AI allows the human expert to appropriately apply their knowledge, improving decision outcomes in cases where the model is likely to perform poorly 167. A previous study also proposes a framework highlighting trust relationships within human-AI teams. It acknowledges the multilevel nature of team trust, which considers individual, dyads, and team levels as a whole 173. Another study points out that the trust level of a team needs to be uniform, as uncooperative members could undermine the team’s ability to reach intended goals 174.
@@
In human-AI decision-making scenarios, people’s awareness of how AI works and its outcomes is critical in building a relationship with the system. AI explanation can support people in justifying their decisions 175. A study on content moderation found that when explanations were provided for why the content was taken down, removal decisions made jointly by humans and AI were perceived as more trustworthy and acceptable than the same decisions made by humans alone 176. Many researchers have proposed using explainable AI (XAI) to enable humans to rely on AI advice appropriately and reach complementary team performance 39. However, some studies suggest that XAI can be associated with a white-box paradox 177, potentially leading to null or detrimental effects.
Previous research finds that AI explanations can increase human acceptance of AI recommendations, regardless of their correctness 178, leading to an over-reliance, threatening the performance of human-AI decision-making. Whereas another study finds that there are scenarios where AI explanations can reduce over-reliance 179 using a cost-benefit framework. The impact of AI explanations on decision-making tasks is contingent upon individuals’ diverse levels of domain expertise 180. Additionally, some studies advocate for the concept of causability 181, which measures whether and to what extent humans can understand a given machine explanation to develop effective human-AI interfaces, particularly in medical AI.
@@
The AI-enhanced collective intelligence approach is increasingly applied across various domains, offering innovative solutions to challenges such as community response to climate change, environment, and sustainability challenges 182, etc. AI serves as a vital decision support tool for policymakers in detecting misinformation 183 and real-time crisis management, exemplified during the COVID-19 pandemic. Its application extends to high-stake areas, including medical diagnosis 184 and criminal justice 185. AI’s integration into collective intelligence efforts enables significant amplification of its impact and scalability 186.
Figure 2: Distribution of AI-enhanced collective intelligence cases by application area based on dataset curated by Supermind Design 187.
Here, we present our analysis utilizing the Supermind Design database 188, which includes over one thousand actual examples of AI-enhanced collective intelligence, with 938 categorized into twelve application areas. Notably, while some cases might not directly involve AI or collective intelligence depending on the exact definitions, the database collectively offers a comprehensive overview. It provides insights into the state-of-the-art in AI-CI applications. As shown in Fig. 2, the majority of applications, approximately twenty percent, are found in the public sector and NGOs, followed by high-tech, media, telecommunications, entertainment, and hospitality. Fewer instances are found in supply chain, real estate, and agriculture. It is important to note that several cases classified under “Public sector, NGO” are private sector initiatives addressing public issues. In the upcoming sections, we will analyze and highlight specific AI-CI examples across these diverse domains. The supplemental information Table S2 provides a summary of the CI and AI aspects of these examples.
@@
Safecity 189 is a platform that crowdsources personal stories of sexual harassment and abuse in public spaces. It collaborates with the AI partner OMDENA 190 to build solutions to help women at risk of sexual abuse by predicting places at a high risk of sexual harassment incidents utilizing convolutional neural networks and LSTM (Long-Short-Term-Memory).
Ushahidi 191 is a non-profit online platform for crowdsourcing information, mapping, and visualizing data to respond to and inform decisions on social issues and crises. It aggregates crowdsourced information and reports from individuals across the globe. Ushahidi is developing new tools to enhance data management and analysis, such as leveraging machine learning techniques to improve the efficiency of processing crisis reports, thus enabling faster response time 192.
While NGOs and the public sector might seem like fertile grounds for AI-CI applications, it is crucial to consider the potential high-risk issues that may arise and their impact during the design and deployment of such AI-CI systems. A notable instance of a data-driven tool for social good that diverged from its original accuracy and effect is the infamous case of Google Flu 193.
Just as AI-enhanced collective intelligence helps address social issues and crises in the public sector and NGOs, it also drives innovation and efficiency in the high-tech industry.
@@
Bluenove 194, a technology and consulting company that pioneers massive collective intelligence, focuses on mobilizing communities on a large scale. Their Assembl 195 platform facilitates open conversations among participants, employing advanced AI technologies such as natural language recognition, semantic analysis, and emotion analysis to categorize content published on the platform automatically.
Figure Eight 196 (formerly known as CrowdFlower 197) was a human-in-the-loop machine learning and AI company acquired by Appen 198 in 2019. It leverages crowdsourced human intelligence to perform tasks such as text transcription or image annotation to train machine learning algorithms, which can be used in various applications.
The recent advancements in high-tech technologies, especially the burgeoning field of LLM software, hold the potential for revolutionary developments. Nevertheless, the high risk associated with errors in automated or AI-generated software 199 necessitates the inclusion of human intelligence in these projects.
@@
Civil War Photo Sleuth 200 combines facial recognition technology and community effort to uncover lost identities in photographs from the American Civil War era. It leverages a dedicated community with a keen interest in the US Civil War and focuses on more accurately tagging and identifying individuals in historical photographs.
Bellingcat 201 is an independent investigative online journalism community specializing in fact-checking and open-source intelligence. It brings together a collective of researchers, investigators, and citizen journalists to publish investigation findings. Utilizing various AI tools in investigative work, Bellingcat obtains insights from massive open-source data, such as satellite imagery, photographs, video recordings, and social media posts.
Facing the emerging challenge of deepfake videos, research indicates that a hybrid system combining human judgment with AI models outperforms either humans or models alone in detecting deepfakes 202, underscoring the value of integrating human perceptual skills with AI technology in addressing the complexities posed by deepfake content.
@@
Swithing to a more industrial environment, in the energy and natural resources sector, collective intelligence combined with AI can play a role in addressing environmental challenges through community-driven data collection and analysis. Litterati 203 is a mobile app and community platform that combats litter by empowering people to take action. The app allows the worldwide community to contribute their observations and findings through photographing and documenting occurrences, creating a crowdsourced database of litter data. It also provides a platform for individuals to connect, share, and support each other’s initiatives. The platform uses AI, such as machine learning and computer vision techniques, to automatically recognize different objects and materials and efficiently organize the collected litter data. Similarly, OpenLitterMap 204 is another example of making a clearer planet using AI-enhanced collective intelligence.
eBird 205 is a mobile app and community platform for biodiversity and supports conservation initiatives to protect bird species and their habitats. It harnesses the power of a global community of birdwatchers and citizen scientists to collectively report bird observations and contribute to a shared knowledge base. eBird uses computer vision techniques to identify bird species from submitted photographs and recordings and employs machine learning predictive models to forecast bird migration patterns and species distributions.
Tackling environmental challenges is crucial, and large-scale, crowd-based AI-enhanced collective intelligence projects play a significant role in this endeavor. Nevertheless, integrating AI with human efforts might dehumanize the work and demotivate human participants 206. This balance between efficiently leveraging AI and maintaining human engagement is a key consideration in such projects.
@@
Education and academia is another area that benefit from AI-enhanced collective intelligence by fostering collaborative learning environments and supporting large-scale research projects. Kialo Edu 207 is an argument mapping and debate site dedicated to facilitating collaborative and critical thinking. Kialo uses natural language processing techniques to analyze arguments, cluster topics, and identify discussion patterns. It allows users to participate in structured debates and discussions, promoting constructive engagement among students. This interaction helps them deeply understand and analyze the core aspects of the topics under discussion.
Zooniverse 208 is a platform for people-powered research where volunteers assist professional researchers. Galaxy Zoo is a prominent project within the Zooniverse platform. This crowdsourced astronomy project invites people to assist in the morphological classification of large numbers of galaxies. It uses AI algorithms to learn patterns and characteristics indicative of different galaxy types by training on large datasets previously classified by citizen scientists. The combination of efforts from both the crowd and AI enables Galaxy Zoo to harness the power of human expertise and the efficiency of AI to scale up the analysis and understanding of galaxies.
A study analyzing the contributor demographics of the Zooniverse project revealed uneven geographical distribution and a gender imbalance, with approximately 30% of the citizen scientists being female 4. This finding highlights concerns about representativeness and diversity in citizen science initiatives, indicating potential biases in the data and insights from such projects. AI can be used to train human contributors 209, to de-bias collective intelligence, and to evaluate and combine human solutions 210.

### v4

@@ Introduction
Much research on collective intelligence has been dedicated to finding a group’s optimum size and structure. Several studies have reported that performance improves with group size through enhanced diversity 17, whereas others suggest the opposite: small groups are more efficient 123. Another study suggests that large teams develop and small teams innovate 124. Larger groups can better utilize diverse perspectives and knowledge to solve complex problems but also tend to experience more coordination problems and communication difficulties 15. The consideration of optimal group size can be related to the dynamics of the group, whether it is static, growing, or diminishing.
In human-AI hybrid teams, group size becomes nuanced with AI integration, especially when AI entities such as LLMs are involved. The countability of AI agents, particularly if multiple agents are derived from a single LLM, raises questions about their distinctiveness and individuality in a team context. Unlike humans, AI agents can simultaneously process multiple tasks, blurring the traditional boundaries of group size. Applying past research on human group dynamics to these hybrid systems requires rethinking the notions of individual contribution, team cohesion, and communication. It is important to consider how AI’s unique capabilities and scale impact these dynamics and whether multiple AI agents from a single model represent distinct entities or a collective intelligence resource.
@@
Research shows that network structure, specifically hierarchy and link intensity, can affect information distribution 125, influencing group performance. Another study observed that the presence of structural holes 126 in leaders’ networks and the adoption of core-periphery and hierarchical structures in groups correlated negatively with their performance 127. Incorporating the temporal dimension, adaptive social networks can significantly impact collective intelligence by allowing the network’s structure to evolve based on feedback from its members 128.
Team size moderates the relationship between network structure and performance 129. Studies have found central members can coordinate with other team members more easily in smaller teams 130, whereas in larger teams, communication challenges may arise 131. A recent study shows a more central connected leader in advice-giving networks has a more positive impact on the performance of larger teams, but this effect is reversed in smaller teams 129.
@@
Teams often outperform individuals primarily due to explicit communication and feedback within the team 132. Understanding the communication network in crowds is crucial for effectively designing and managing crowdsourcing tasks 133. Various studies highlight that centralized and decentralized communication patterns can effectively promote team performance, contingent on the nature of the task and the team’s composition 134, 135. In particular, teams handling complex tasks tend to be more productive with decentralized communication networks 136. Conversely, the centralization of communication around socially dominant or less reflective individuals can negatively impact the utilization of expertise and overall team performance 137. Communication patterns, such as equal participation and turn-taking of team members, can positively affect collective intelligence 15.
The integration of AI in human groups can influence the conversation dynamics in different ways, depending on the role of AI. AI chat interventions can improve online political conversations at scale 138, reducing the chance of conflicts. While AI machines might struggle with capturing the subtle and ineffable social expressions that make up the dynamics of human groups, this can be advantageous. In certain contexts, AI can be social catalysts 139 to promote communication where human capacity is limited. Striking the right balance in AI’s role is key to maintaining the natural dynamics of human interaction.
@@
Cognitive processes are foundations for the emergence of collective intelligence 6, encompassing knowledge acquisition, information processing, problem-solving, decision-making, language, perception, memory, attention, and reasoning. Human cognition processes evolve and interact with each other during various tasks and experiences. Humans can represent the mental states of others, including desires, beliefs, and intentions 140, known as the Theory of Mind (ToM). Past research, inspired by ToM, showed that team members’ ability to assess others’ mental state is positively associated with team performance both in face-to-face settings 15 and online 141. Regarding collective understanding within a group, the Shared Mental Models (SMM) theory is helpful for understanding, predicting, and improving performance in human teams 142.
Applying theories such as SMM to human-AI teams is a natural progression in research. These models can enhance team performance by ensuring that human and AI members have a similar understanding of tasks and each other 142, facilitating better prediction of needs and behaviors. Developing ToM in AI also improves coordination and adaptability in human-AI teams 143. This involves AI’s capability to anticipate and respond to new information and behaviors. Moreover, recent studies indicate that incorporating “hot” cognition, influenced by emotional states in AI, can further improve human-machine interactions 144. Understanding the transactive systems 19 of memory, attention, and reasoning within these teams is key to grasping the members’ knowledge and task preferences.
@@
Assistant-type AI systems generally function as technical tools with limited autonomy, designed to complement or augment human abilities and enhance efficiency in task performance. For example, language translation AI can assist in translating text or spoken language, facilitating communication across language barriers. LLMs in education can assist students’ learning by adapting to different roles based on the prompts provided by the students 145. Administrative AI assistants can help schedule meetings, manage emails, and organize tasks. Smart home AI assistants such as Alexa and Siri can help manage smart home devices, play music, provide weather updates, answer questions, and assist with daily routines. AI assistants are technical tools that help humans coordinate tasks efficiently, optimize their decision-making, and personalize their experiences.
@@
AI can function as teammates, working alongside humans with complementary abilities. AI teammates are already used in real-world settings; most employees using AI already see it as a coworker. In healthcare, radiologists and AI work together to diagnose from the pictures of the patients. In creative industries, AI collaborators in music, art, and literature, such as Google’s Magenta project, collaborate with musicians and artists to create new compositions or artworks using AI algorithms. Furthermore, AI robots can serve as teammates in various workplaces, a topic we will discuss in more detail later.
@@
As a coach, AI can provide guidance, support, and personalized feedback to individuals seeking to improve their skills and achieve their goals. In team settings, one perspective highlights the AI coach’s ability to provide a comprehensive, global view of the team’s environment, guiding players who have only partial views 146. This approach involves the AI coach strategizing and coordinating, distributing tailored strategies to each team member based on their unique perspectives and roles. Another dimension of AI coaching focuses on assessing and improving teamwork by closely monitoring team members during collaborative tasks 147. Here, the AI coach actively intervenes with timely suggestions based on the team members’ inferred mental model misalignment 147.
@@
AI has been used in working environments with supervision or managerial roles, such as making decisions of hiring, promotion, and reassigning tasks 148. An AI manager may mitigate human biases 149 in hiring and promotions, contributing to a more diverse workforce, which is a key aspect of collective intelligence. Additionally, by analyzing individual strengths and weaknesses and efficiently allocating tasks, an AI manager can optimize team dynamics and workflows, improving overall group performance.
@@
With the advancement of AI in areas such as natural language processing, object recognition, and creative idea generation, there’s a trend to integrate AI with robotics, creating physical entities in the real world. Embodied AI 150 refers to intelligent agents interacting with their environment through a physical body. These AI-enabled machines, embodied in robotic arms or autonomous vehicles, can augment the capabilities of human workers. They are equipped with advanced sensors, motors, and actuators, enabling recognition of people and objects and safe collaboration with humans in diverse settings, including factories, warehouses, and laboratories. Furthermore, advanced AI-robotics integration allows robots to recognize and respond to human speech and gestures, impacting conversational dynamics within mixed human-robot environments. Research has shown that a vulnerable robot’s social behavior positively shapes the conversation dynamics among human participants in a human-robot group 151. In scenarios where conflicts emerge within a human group or in groups of strangers, a humorous robot can potentially serve as an ice breaker to intervene and resolve the awkward situation 152, demonstrating the potential of robots to influence the nature of human-human interactions notably.
@@
In addition to investigating how humans perceive individual AI, previous research suggests the necessity of additional work to explore human perceptions of AI in collaborative team environments 153. For human-AI teams to succeed, human team members must be receptive to their new AI counterparts 154. People’s pre-existing attitudes toward AI were found to be significantly related to their willingness to be involved in human-AI teams 155. Research finds that two aspects of social perception, warmth, and competence, are critical predictors of human receptivity to a new AI teammate. Psychological acceptance is positively related to perceived human-AI team viability 154. Regarding human perception of AI behaviors, one study finds that people cooperate less with benevolent AI agents than with benevolent humans 156.
Humans tend to be more accepting of AI when it makes mistakes. Some studies find that users have a more favorable impression of the imperfect robot than the perfect robot when the robot behaves adequately after making mistakes 157. When errors occur in human-machine shared-control vehicles, the blame assigned to the machine is reduced if both drivers make mistakes 158. Research has found that errors occasionally performed by a humanoid robot can increase its perceived human likeness and likability 159. A recent study shows humans accept their AI teammate’s decision less often when they are deceived about the identity of the AI as another human 160. The nature of the task requested by the robot, e.g., whether its effects are revocable as opposed to irrevocable, significantly impacts participants’ willingness to follow its instructions 161. When taking AI’s advice, humans must know the AI system’s error boundary and decide when to accept or override AI’s recommendation 162.
@@
In AI-assisted decision-making, where the individual strengths of humans and AI optimize the joint decision outcome, the key to success is appropriately calibrating human trust 163. Trust between humans and AI plays a pivotal role in realizing collective intelligence 164. Previous research highlights the necessity of understanding factors that enable or hinder the formation, maintenance, and repair of trust in human-AI collaborations 164. Key factors influencing this trust include the perceived competence, benevolence, and integrity 165 of the AI systems, mirroring the trust dynamics in human relationships.
The level of anthropomorphism of machines can be a predictor of the trust level of the human participants 166. Studies indicate that while anthropomorphic features in AI agents initially create positive impressions, this effect is often short-lived 164. Trust tends to decline, especially when the agent’s capabilities are not clearly presented and its performance fails to meet users’ expectations 167.
Trust in a teammate may evolve over time irrespective of team performance 168. Knowing when to trust the AI allows the human expert to appropriately apply their knowledge, improving decision outcomes in cases where the model is likely to perform poorly 163. A previous study also proposes a framework highlighting trust relationships within human-AI teams. It acknowledges the multilevel nature of team trust, which considers individual, dyads, and team levels as a whole 169. Another study points out that the trust level of a team needs to be uniform, as uncooperative members could undermine the team’s ability to reach intended goals 170.
@@
In human-AI decision-making scenarios, people’s awareness of how AI works and its outcomes is critical in building a relationship with the system. AI explanation can support people in justifying their decisions 171. A study on content moderation found that when explanations were provided for why the content was taken down, removal decisions made jointly by humans and AI were perceived as more trustworthy and acceptable than the same decisions made by humans alone 172. Many researchers have proposed using explainable AI (XAI) to enable humans to rely on AI advice appropriately and reach complementary team performance 37. However, some studies suggest that XAI can be associated with a white-box paradox 173, potentially leading to null or detrimental effects.
Previous research finds that AI explanations can increase human acceptance of AI recommendations, regardless of their correctness 174, leading to an over-reliance, threatening the performance of human-AI decision-making. Whereas another study finds that there are scenarios where AI explanations can reduce over-reliance 175 using a cost-benefit framework. The impact of AI explanations on decision-making tasks is contingent upon individuals’ diverse levels of domain expertise 176. Additionally, some studies advocate for the concept of causability 177, which measures whether and to what extent humans can understand a given machine explanation to develop effective human-AI interfaces, particularly in medical AI.
@@
The AI-enhanced collective intelligence approach is increasingly applied across various domains, offering innovative solutions to challenges such as community response to climate change, environment, and sustainability challenges 178, etc. AI serves as a vital decision support tool for policymakers in detecting misinformation 179 and real-time crisis management, exemplified during the COVID-19 pandemic. Its application extends to high-stake areas, including medical diagnosis 180 and criminal justice 181. AI’s integration into collective intelligence efforts enables significant amplification of its impact and scalability 182.
Here, we present our analysis utilizing the Supermind Design database 183, which includes over one thousand actual examples of AI-enhanced collective intelligence, with 938 categorized into twelve application areas. Notably, while some cases might not directly involve AI or collective intelligence depending on the exact definitions, the database collectively offers a comprehensive overview. It provides insights into the state-of-the-art in AI-CI applications. As shown in Fig. 2, the majority of applications, approximately twenty percent, are found in the public sector and NGOs, followed by high-tech, media, telecommunications, entertainment, and hospitality. Fewer instances are found in supply chain, real estate, and agriculture. It is important to note that several cases classified under “Public sector, NGO” are private sector initiatives addressing public issues. In the upcoming sections, we will analyze and highlight specific AI-CI examples across these diverse domains. The supplemental information Table S2 provides a summary of the CI and AI aspects of these examples.
@@
Safecity (safecity.in) is a platform that crowdsources personal stories of sexual harassment and abuse in public spaces. It collaborates with the AI partner OMDENA to build solutions to help women at risk of sexual abuse by predicting places at a high risk of sexual harassment incidents utilizing convolutional neural networks and LSTM (Long-Short-Term-Memory).
Ushahidi (www.ushahidi.com) is a non-profit online platform for crowdsourcing information, mapping, and visualizing data to respond to and inform decisions on social issues and crises. It aggregates crowdsourced information and reports from individuals across the globe. Ushahidi is developing new tools to enhance data management and analysis, such as leveraging machine learning techniques to improve the efficiency of processing crisis reports, thus enabling faster response time 184.
While NGOs and the public sector might seem like fertile grounds for AI-CI applications, it is crucial to consider the potential high-risk issues that may arise and their impact during the design and deployment of such AI-CI systems. A notable instance of a data-driven tool for social good that diverged from its original accuracy and effect is the infamous case of Google Flu 185.
@@
Just as AI-enhanced collective intelligence helps address social issues and crises in the public sector and NGOs, it also drives innovation and efficiency in the high-tech industry. Bluenove (bluenove.com), a technology and consulting company that pioneers massive collective intelligence, focuses on mobilizing communities on a large scale. Their Assembl platform (bluenove.com/en/offers/assembl) facilitates open conversations among participants, employing advanced AI technologies such as natural language recognition, semantic analysis, and emotion analysis to categorize content published on the platform automatically.
Figure Eight (formerly known as CrowdFlower was a human-in-the-loop machine learning and AI company acquired by Appen (appen.com) in 2019. It leverages crowdsourced human intelligence to perform tasks such as text transcription or image annotation to train machine learning algorithms, which can be used in various applications.
The recent advancements in high-tech technologies, especially the burgeoning field of LLM software, hold the potential for revolutionary developments. Nevertheless, the high risk associated with errors in automated or AI-generated software 186 necessitates the inclusion of human intelligence in these projects.
@@
Civil War Photo Sleuth (civilwarphotosleuth.com) combines facial recognition technology and community effort to uncover lost identities in photographs from the American Civil War era. It leverages a dedicated community with a keen interest in the US Civil War and focuses on more accurately tagging and identifying individuals in historical photographs.
Bellingcat (bellingcat.com) is an independent investigative online journalism community specializing in fact-checking and open-source intelligence. It brings together a collective of researchers, investigators, and citizen journalists to publish investigation findings. Utilizing various AI tools in investigative work, Bellingcat obtains insights from massive open-source data, such as satellite imagery, photographs, video recordings, and social media posts.
Facing the emerging challenge of deepfake videos, research indicates that a hybrid system combining human judgment with AI models outperforms either humans or models alone in detecting deepfakes 187, underscoring the value of integrating human perceptual skills with AI technology in addressing the complexities posed by deepfake content.
@@
Swithing to a more industrial environment, in the energy and natural resources sector, collective intelligence combined with AI can play a role in addressing environmental challenges through community-driven data collection and analysis. Litterati 188 is a mobile app and community platform that combats litter by empowering people to take action. The app allows the worldwide community to contribute their observations and findings through photographing and documenting occurrences, creating a crowdsourced database of litter data. It also provides a platform for individuals to connect, share, and support each other’s initiatives. The platform uses AI, such as machine learning and computer vision techniques, to automatically recognize different objects and materials and efficiently organize the collected litter data. Similarly, OpenLitterMap (openlittermap.com) is another example of making a clearer planet using AI-enhanced collective intelligence.
eBird 189 is a mobile app and community platform for biodiversity and supports conservation initiatives to protect bird species and their habitats. It harnesses the power of a global community of birdwatchers and citizen scientists to collectively report bird observations and contribute to a shared knowledge base. eBird uses computer vision techniques to identify bird species from submitted photographs and recordings and employs machine learning predictive models to forecast bird migration patterns and species distributions.
Tackling environmental challenges is crucial, and large-scale, crowd-based AI-enhanced collective intelligence projects play a significant role in this endeavor. Nevertheless, integrating AI with human efforts might dehumanize the work and demotivate human participants 190. This balance between efficiently leveraging AI and maintaining human engagement is a key consideration in such projects.
@@
Education and academia are other areas that benefit from AI-enhanced collective intelligence by fostering collaborative learning environments and supporting large-scale research projects. Kialo Edu (kialo-edu.com) is an argument mapping and debate site dedicated to facilitating collaborative and critical thinking. Kialo uses natural language processing techniques to analyze arguments, cluster topics, and identify discussion patterns. It allows users to participate in structured debates and discussions, promoting constructive engagement among students. This interaction helps them deeply understand and analyze the core aspects of the topics under discussion.
Zooniverse (zooniverse.org) is a platform for people-powered research where volunteers assist professional researchers. Galaxy Zoo is a prominent project within the Zooniverse platform. This crowdsourced astronomy project invites people to assist in the morphological classification of large numbers of galaxies. It uses AI algorithms to learn patterns and characteristics indicative of different galaxy types by training on large datasets previously classified by citizen scientists. The combination of efforts from both the crowd and AI enables Galaxy Zoo to harness the power of human expertise and the efficiency of AI to scale up the analysis and understanding of galaxies.
A study analyzing the contributor demographics of the Zooniverse project revealed uneven geographical distribution and a gender imbalance, with approximately 30% of the citizen scientists being female 4. This finding highlights concerns about representativeness and diversity in citizen science initiatives, indicating potential biases in the data and insights from such projects. AI can be used to train human contributors 191, to de-bias collective intelligence, and to evaluate and combine human solutions 192.